Solve each system by the substitution method
a) 2x - y = 4
2x - y = 3
b) x + 3y = 2
-x + y = 1

Answer by stanbon(75887)

Solve each system by the substitution method
a)
2x - y = 4
2x - y = 3
------
1st equation: y = 2x-4
---
Substitute for y in 2nd and solve for "x":
2x - (2x-4) = 3
2x-2x+4 = 3
4 = 3
---
Contradiction: Therefore No solution
==========================================
b)
x + 3y = 2
-x + y = 1
---
1st equation: x = -3y+2
---
Substitute for "x" and solve for "y":
-(-3y+2)+y = 1
3y-2 + y = 1
4y = 3
y = 3/4
---------
Since x = -3y+2, x = -3(3/4)+2 = -9/4+8/4 = -1/4
======================================================
Cheers,
Stan H.